OpenStudy (phebe):

A transformer consists of 250 primary windings and 725 secondary windings. If the emf of the primary coil is 20.0 V, what is the emf across the secondary coil? A- 29.0 V B- 18.1 V C- 58.0 V D- 34.5 V

OpenStudy (phebe):

i think its gonna be 58.0

OpenStudy (phebe):

i think you multiply the emf not sure

OpenStudy (phebe):

theres a formula for this right?

OpenStudy (mayankdevnani):

yeah ! \[\large \bf \frac{secondary~windings}{primary~widnings}=\frac{EMF~secondary}{EMF~primary}\]

OpenStudy (mayankdevnani):

just plug all the values and get your answer

OpenStudy (phebe):

okay thanks!

OpenStudy (phebe):

let me see what i get

OpenStudy (mayankdevnani):

sure!

OpenStudy (phebe):

725/250 right?

OpenStudy (mayankdevnani):

yep! this is your left hand side fraction

OpenStudy (phebe):

okay so its 2.9

OpenStudy (phebe):

So is it A?

OpenStudy (mayankdevnani):

right !

OpenStudy (phebe):

YES! THANK YOUUU!!!! <3

OpenStudy (mayankdevnani):

\[\large \bf \frac{725}{250}=\frac{x}{20}\]

OpenStudy (mayankdevnani):

find x ?

OpenStudy (phebe):

SO DIVIDE 250 and 20?

OpenStudy (mayankdevnani):

\[\large \bf 2.9=\frac{x}{20}\]

OpenStudy (mayankdevnani):

multiply both sides by 20,we get \[\large \bf 2.9 \times 20=x\]

OpenStudy (mayankdevnani):

now solve for x

OpenStudy (phebe):

so its 58 That means my original answer was always right thanks!

OpenStudy (mayankdevnani):

np !